OpenStudy (anonymous):

Suppose an object is dropped from a height h0 above the ground. Then its height after t seconds is given by h = −16t ^2 + h0, where h is measured in feet. Use this information to solve the problem. A ball is dropped from the top of a building 76 ft tall. (Round your answers to three decimal places.) (a) How long will it take to fall half the distance to ground level? t = ____ sec (b) How long will it take to fall to ground level? t =____ sec

OpenStudy (anonymous):

\[h = h_o/2 = -16*t^2+h_o \]\ thus [(h_0/32)^(1/2)=t\] for part a

OpenStudy (anonymous):

a. t = (h_0/32)^(1/2) b. t = (h_0/16)^(1/2)

OpenStudy (anonymous):

rearranging equation gives t=sqr ( [-h +h0] / 16 ) ergo (a) t = 1.541 (b) t = 2.179

OpenStudy (anonymous):

a. t = (h_0/32)^(1/2) b. t = (h_0/16)^(1/2)